OpenStudy (anonymous):

Determine the end behavior of each function.As x aproaches postive and negative infinty (fx) = 9x^2 − 11x + 2/8x^2 − 2

OpenStudy (anonymous):

I think you should divide by \(x^2\) and then see what happens when you substitute infinity in for x.

OpenStudy (anonymous):

I dont see how i could divide by x^2

OpenStudy (anonymous):

\[\frac{9x^2}{x^2}-\frac{11x}{x^2}+\frac{2/8x^2}{x^2}-\frac{2}{x^2} \]